OpenStudy (anonymous):

A simple syrup used for ice cream toppings requires 2 cups of sugar and 2/3 cup of boiling water. How many cups of sugar are required for 2 cups of boiling water?

OpenStudy (anonymous):

1 syrup is contains 2 sugar + 2/3 water if n cups of boiling of water is 2, so ( times all above by 3) 3*(1 syrup) = 3*(2 sugar) + 3*(2/3 water) 3 syrup = 6 sugar + 2 water.., so, the cups of sugar are required for 2 cups of boiling water is 6 for 3 syrup @rmaker408

OpenStudy (anonymous):

thanks!! @gerryliyana